### v4

# Supplementary for Online Price Competition under Generalized Linear Demands

Abstract
We study sequential price competition among $N$ sellers, each influenced by the pricing decisions of their rivals. Specifically, the demand function for each seller $i$ follows the single index model $\lambda_{i}(\mathbf{p})=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle)$, with known increasing link $\mu_{i}$ and unknown parameter $\boldsymbol{\theta}_{i,0}$, where the vector $\mathbf{p}$ denotes the vector of prices offered by all the sellers simultaneously at a given instant. Each seller observes only their own realized demand – unobservable to competitors – and the prices set by rivals. Our framework generalizes existing approaches that focus solely on linear demand models. We propose a novel decentralized policy, PML-GLUCB, that combines penalized MLE with an upper-confidence pricing rule, removing the need for front-loaded exploration phases across sellers — which is integral to previous linear models — and accommodating both binary and real-valued demand observations. Relative to a dynamic benchmark policy, each seller achieves $O(N^{2}\sqrt{T}\log(T))$ regret, which essentially matches the optimal rate known in the linear setting. A significant technical contribution of our work is the development of a variant of the elliptical potential lemma — typically applied in single-agent systems — adapted to our competitive multi-agent environment.
Machine Learning, ICML
1 Introduction
A pricing strategy aims to determine prices that maximize current profits. Much of the learning-to-price literature, however, has traditionally focused on monopolistic settings (Fan et al., 2024; Javanmard and Nazerzadeh, 2019; Golrezaei et al., 2019; Bracale et al., 2025b), effectively treating demand as independent of competitors’ actions. In many real-world markets, this assumption is unrealistic. Markets typically feature multiple sellers offering products that differ in quality and price, each seeking to maximize its own revenue. In such environments, a seller’s profit is inherently affected by the pricing decisions of its competitors. As a result, firms must continuously revise their prices to anticipate and respond to rivals’ moves. Motivated by these considerations, recent work has moved towards the study of pricing policies in multi-seller environments, explicitly modeling strategic interactions among competing firms (Li et al., 2024; Bracale et al., 2025a; Yang et al., 2024; Goyal et al., 2023; Zhu et al., 2025). In such settings, each seller seeks to maximize its own revenue, which inherently depends on the pricing policies adopted by its competitors. This interdependence naturally leads to the concept of a Nash equilibrium: a profile of prices at which no seller can improve its payoff through a unilateral deviation, given the strategies of the others. Building on this framework, existing works (Li et al., 2024; Bracale et al., 2025a) propose sequential pricing mechanisms designed to learn equilibrium behavior over time, and establish sublinear regret guarantees relative to the equilibrium benchmark [^1] In this paper, we use the terms sublinear regret and no regret interchangeably to indicate that the cumulative regret grows sub-linearly with time, that is, $\mathrm{Reg}(T)=o(T)$ as $T\to\infty$. and converge to a Nash equilibrium (NE) (Li et al., 2024; Bracale et al., 2025a). Yet two practical issues remain:
  1. 1.
    Existing non-monopolistic approaches (Li et al., 2024; Bracale et al., 2025a) typically decouple exploration” from optimization” and rely on an explicit, front-loaded exploration phase with an i.i.d. pricing policy. However, as noted by Jin et al. (2021), classical results such as Garivier et al. (2016) show that the explore-then-commit strategy can be asymptotically suboptimal as the horizon grows, and can therefore be dominated by fully sequential policies such as Upper Confidence Bound (UCB). Moreover, in some dynamic-pricing implementations, the required length of the exploration phase depends on unknown problem-dependent constants (Bracale et al., 2025a), which complicates deployment in practice.
  2. 2.
    The demand is often forced to be linear (or narrowly parameterized) (Bertsimas and Perakis, 2006; Guda and Subramanian, 2026; Li et al., 2024; Goyal et al., 2023), limiting the ability to capture nonlinear demand patterns commonly observed in real markets (Gallego et al., 2006; Wan et al., 2022, 2023).
To address these limitations, we adopt a generalized linear (single-index) demand framework, which accommodates a broader range of nonlinear demand responses and naturally covers both binary and continuous demand outcomes, and propose a pricing policy that operates without a dedicated initial i.i.d. pricing-design exploration phase. We prove that each seller’s regret remains on the order of $\widetilde{O}(\sqrt{T})$ against a dynamic benchmark, where $\widetilde{O}$ suppresses logarithmic factors.
1.1 Related Literature
Demand learning in competitive environments.
Over the past two decades, there has been substantial progress in demand learning under competition. Early work by Kirman (1975) analyzed symmetric two-seller competition with linear demand. Asymmetry and noise were incorporated by Cooper et al. (2015), who let each seller learn independently (effectively as a monopolist) and established equilibrium under known price sensitivities. In contrast, Li et al. (2024) studied dynamic pricing with unknown price sensitivities under a linear demand model and achieved the optimal $\widetilde{O}(\sqrt{T})$ regret rate. In their setup, seller $i$’s expected demand is
$\lambda_{i}(\mathbf{p})=\alpha_{i,0}-\beta_{i,0}p_{i}+\sum_{j\in[N]\setminus\{i\}}\gamma_{ij,0}p_{j},$
where $N]=\{1,2,\dots,N\}$ indexes the sellers and the parameters $(\alpha_{i,0},\beta_{i,0},\gamma_{ij,0})$ are unknown with $\beta_{i,0}>0$. The condition $\beta_{i,0}>0$ captures the standard economic assumption that the demand of seller $i$ decreases as their own price $p_{i}$ increases. However, nonlinear demand models are often more realistic and have been explored in several applications (Gallego et al., 2006; Wan et al., 2022, 2023). For instance, Gallego et al. (2006) considered the mean demand function $\lambda_{i}(\mathbf{p})=a_{i}\left(p_{i}\right)/(\sum_{j}a_{j}\left(p_{j}\right)+\kappa)$, for some known increasing functions $a_{i}$ and real value $\kappa\in(0,1]$. They establish the existence of a NE, but they do not propose an online learning procedure.
In contrast to prior studies (Guda and Subramanian, 2026; Li et al., 2024; Kachani et al., 2007; Gallego et al., 2006), we analyze a more general monotone single-index expected-demand model:
$\lambda_{i}(\mathbf{p})=\mu_{i}(-\beta_{i,0}p_{i}+\langle\boldsymbol{\gamma}_{i,0},\mathbf{p}_{-i}\rangle)=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle),$ (1)
for each $i\in[N]$, where $\mathbf{p}_{-i}=(p_{j})_{j\in[N]\setminus\{i\}}$, $\mathbf{p}=(p_{j})_{j\in[N]}$ and $\boldsymbol{\theta}_{i,0}$ is a vector of dimension $N$ with $i$-th entry equals to $\beta_{i,0}$ (where $\beta_{i,0}>0$) and the rest of the entries are the values of $\boldsymbol{\gamma}_{i,0}$ ordered. The link function $\mu_{i}$ is assumed to be increasing, capturing the same economic intuition that demand decreases with the seller’s own price $p_{i}$. Here $\mu_{i}$ is known to seller $i$ (but unknown to competitors) and the parameter $\boldsymbol{\theta}_{i,0}$ is unknown (to everyone). This model generalizes the linear demand model by Li et al. (2024); specifically, $\mu_{i}(u)=u+\alpha_{i,0}$ recovers their formulation.
Multi-Agent Reinforcement Learning (MARL).
MARL is a broad framework for sequential decision-making problems involving multiple strategic agents that at each instant select actions based on their local observations, receive feedback in the form of rewards (which depend on personal and competitors’ actions), and then update their policy to maximize long-term returns. The ultimate objective is to reach a stable solution concept, such as an NE, where no agent can unilaterally improve its expected return.
A key modeling assumption in MARL algorithms is full or partial observability of other agents’ rewards or payoffs (Hu and Wellman, 2003; Zhang et al., 2021; Yang and Wang, 2020). For instance, the canonical formulation of Hu and Wellman (2003) assumes that each agent can observe the immediate reward of all players and use this to update joint policy. Similarly, Zhang et al. (2021) and Yang and Wang (2020) rely on either joint reward observability or access to sufficient statistics (e.g., payoff signals or utility vectors) that allow agents to infer the equilibrium structure. This assumption is critical: observing other agents’ realized rewards enables an agent to estimate the full payoff landscape and to learn best-response dynamics with respect to the joint reward function.
Our setting fundamentally differs from this MARL paradigm in both information structure and observability. In the sequential pricing game we study, each seller observes only their own reward (or demand value) and rivals’ posted prices – never competitors’ demands or revenues. Because demand functions are private and unobservable, a seller cannot recover rivals’ payoff functions or utility gradients, so classical MARL techniques are not directly applicable. Nevertheless, we design a policy whose dynamics need not converge to a Nash equilibrium, yet whose individual regret satisfies $\widetilde{O}(\sqrt{T})$. This makes our approach better aligned with real-world competitive pricing, where revenue data is proprietary, and competitors’ demand is hidden.
Connection to contextual bandits – and a timing asymmetry.
One may view $\mathbf{p}_{-i}^{(t)}$ at any instant $t$ as the “context” for seller $i$, linking our problem to contextual bandits (Chu et al., 2011; Agrawal and Goyal, 2013; Chakraborty et al., 2023; Kim et al., 2023) and to the nonlinear Cournot structure (Cournot, 1838; Saloner, 1987). That said, our model departs from the standard contextual bandit framework in two key ways. First, in the bandit literature, the context is typically exogenous — sampled independently of the agent’s actions — whereas here the “context” $\mathbf{p}_{-i}^{(t)}$ evolves as a function of past prices (and thus past demand noise realizations). Second, classical contextual bandits assume the context is observed before the action is chosen; in our market, sellers post prices simultaneously, so seller $i$ observes $\mathbf{p}_{-i}^{(t)}$ only after selecting $p_{i}^{(t)}$. This delayed observation prevents a direct application of standard contextual bandit analyses, in particular arguments based on the elliptical potential lemma (Carpentier et al., 2020), which is central to many regret proofs. We address this timing mismatch by adapting the analysis to our simultaneous-move setting, deriving an appropriate elliptical-type bound (see Lemma C.2) that yields regret control despite the delayed “context”.
1.2 Contributions
We summarize our main contributions.
[…]
(2) Generalize linear demand with private feedback. We study a competitive demand-learning model of the form $\lambda_{i}(\mathbf{p})=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle),$ where $\mu_{i}$ is a known increasing link and $\boldsymbol{\theta}_{i,0}$ is unknown. This strictly generalizes the linear models (Li et al., 2024; Guda and Subramanian, 2026) and, in contrast to prior works, naturally covers both (i) binary outcomes (purchase / no-purchase) (Goyal et al., 2023; Golrezaei et al., 2019; Javanmard, 2017; Javanmard and Nazerzadeh, 2019; Fan et al., 2024; Bracale et al., 2025b) and (ii) continuous demand observations (Bertsimas and Perakis, 2006; Li et al., 2024; Bracale et al., 2025a) within the same analysis. Crucially, our information structure matches real markets: each seller observes only their own realized demand/revenue and competitors’ posted prices, not competitors’ outcomes, ruling out standard multi-agent RL assumptions (as noted above in Section 1.1) and requiring new proof techniques.
[…]
1.3 General Notation
[…]
1.4 Organization of the paper
The remainder of the paper is organized as follows. Section 2 introduces the problem formulation and model assumptions, and formally defines regret and Nash equilibrium. Section 3 presents our proposed methods, including the penalized maximum likelihood estimator and the UCB-based pricing policy. Section 4 derives concentration bounds and establishes our main regret guarantees, outlining the key steps of the proofs. Section 5 concludes the paper by discussing limitations and future research directions. Additional technical details and complete proofs are deferred to the supplementary material.
2 Problem Setting
We consider a pricing problem faced by $N$ sellers that sell a single product with unlimited inventory sequentially over a time horizon $T$, where $T$ is known to the sellers.
Assumption 2.1 (Bounded prices).
For all $t\in[T]$, the $i$-th seller’s price is denoted as $p_{i}^{(t)}\in\mathcal{P}_{i}\triangleq[\underline{p_{i}},\bar{p}_{i}]$, with price bounds $0<\underline{p_{i}}<\bar{p}_{i}<+\infty$. Every seller knows their own price domain $\mathcal{P}_{i}$ but not the competing ones $\{\mathcal{P}_{j}\}_{j\in[N]\setminus\{i\}}$. However, only $B_{p}=(\sum_{i\in[N]}\bar{p}_{i}^{2})^{1/2}$ is known to all the sellers.
Let $\mathbf{p}_{-i}^{(t)}\triangleq(p_{j}^{(t)})_{j\in[N]\setminus\{i\}}$ denote the $i$-th seller’s competitors’ prices at time $t$ and $\mathbf{p}^{(t)}\triangleq(p_{j}^{(t)})_{j\in[N]}$ denote the joint prices. Only for the purpose of the analysis we define $\mathcal{P}\triangleq\prod_{i\in[N]}[\underline{p_{i}},\bar{p}_{i}]$ which denotes the support of joint prices and $\mathcal{P}_{-i}=\prod_{j\in[N]\setminus\{i\}}\mathcal{P}_{j}$ the support of the competitors’ prices. The general dynamic is as follows: at time $t=0$, each seller $i$ selects any initial price $p_{i}^{(0)}\in\mathcal{P}_{i}$ and everyone observes their own realized demand $y_{i}^{(0)}$ according to (2). The price vector $\mathbf{p}^{(0)}$ is made public. Let $\mathcal{H}^{(0)}_{i}=\{(\mathbf{p}^{(0)},y_{i}^{(0)})\}$ be the data collected by seller $i$. For each sales round $t=1,\ldots,T$:
  1. 1.
    Each seller $i$ sets $p_{i}^{(t)}$ based on their history $\mathcal{H}_{i}^{(t-1)}$.
  2. 2.
    Each seller $i$ observes their realized demand $y_{i}^{(t)}$, sampled by nature according to the model in (2). The vector of prices $\mathbf{p}^{(t)}$ is made public, and then each seller updates their history: $\mathcal{H}_{i}^{(t)}=\mathcal{H}_{i}^{(t-1)}\cup\{(\mathbf{p}^{(t)},y_{i}^{(t)})\}$.
  3. 3.
    Each seller $i$ observes their revenue $r_{i}=p_{i}^{(t)}y_{i}^{(t)}$.
We emphasize that sellers can observe the historical prices of competitors, but do not know the demand of competitors. We assume that $y_{i}^{(t)}$, given the price vector $\mathbf{p}^{(t)}$, is sampled by nature according to a canonical exponential family w.r.t. a reference measure $\nu_{i}$:
$\frac{d\mathbb{P}_{\boldsymbol{\theta}_{i,0}}(y_{i}|\mathbf{p})}{d\nu_{i}(y_{i})}=\exp\left\{y_{i}\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle-b_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle)+c_{i}(y_{i})\right\},$ (2)
where $c_{i}(\cdot)$ is a real-valued function and $b_{i}(\cdot)$ is assumed to be twice continuously differentiable. A random variable $y_{i}$ with the above density satisfies $\mathbb{E}(y_{i}|\mathbf{p})=b^{\prime}_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle)$ and $\textnormal{var}(y_{i}|\mathbf{p})=b^{\prime\prime}_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle)$, showing that $b_{i}(\cdot)$ is strictly convex. The inverse link function $\mu_{i}:=b^{\prime}_{i}$ is consequently strictly increasing. Thus, we can write
$\lambda_{i}(\mathbf{p}^{(t)})$ $=\mathbb{E}[y_{i}^{(t)}|\mathbf{p}^{(t)}]$
$=\mu_{i}(-\beta_{i,0}p_{i}^{(t)}+\langle\boldsymbol{\gamma}_{i,0},\mathbf{p}_{-i}^{(t)}\rangle)=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}^{(t)}\rangle),$
where $\mu_{i}$ is known to the seller $i$ (but not to competitors $j\in[N]\setminus\{i\}$) and $\boldsymbol{\theta}_{i,0}$ is an $N$ dimensional vector unknown to everyone with $i$-th entry equals to $\beta_{i,0}$ and the rest of the entries are the values of $\boldsymbol{\gamma}_{i,0}$ ordered. The parameters $\beta_{i,0}$ and $\boldsymbol{\gamma}_{i,0}$ are commonly referred to as the $i$-th seller’s own-price sensitivity and the competitors’ price sensitivities, respectively. The demands $\{y_{j}^{(t)}\}_{j\in[N]}$ within a period $t$ can be correlated across sellers conditional on $\mathbf{p}^{(t)}$. The model in (2) generates samples $y_{i}^{(t)}$ that reflect not only each seller’s own price $p_{i}^{(t)}$ but also the strategies of the competitors $\mathbf{p}^{(t)}_{-i}=(p_{j}^{(t)})_{j\in[N],j\neq i}$. The exponential model in (2) allows us to generalize linear models and also to accommodate both binary and continuous outcomes: when binary, it might represent the indicator of whether seller $i$ sold the item or not at price $p_{i}^{(t)}$; when continuous, it can be seen as observed mean demand value at price $\mathbf{p}^{(t)}$ plus a mean-zero error. We present two relevant examples.
Example 2.2 (Binary Response Model).
We consider a generalization of the monopolistic binary response model considered by Fan et al. (2024); Bracale et al. (2025b) (and corresponding references). Each customer arrives at time $t$ and, given the price vector posted by sellers $\mathbf{p}^{(t)}$, samples a valuation for each seller $i$
$v_{i}^{(t)}=\langle\widetilde{\boldsymbol{\theta}}_{i,0},\mathbf{p}^{(t)}\rangle+\delta_{i}^{(t)},$
where $\widetilde{\boldsymbol{\theta}}_{i,0}$ is an unknown parameter and $\delta_{i}^{(t)}$ are i.i.d. across $t$ with c.d.f. $F_{i}$. Here $\delta_{i}^{(t)}$ are independent across $t$ but might be correlated across $i$ conditional on $\mathbf{p}^{(t)}$. Let $p_{i}^{(t)}$ be the price posted by seller $i$. A purchase from $i$ happens if $y_{i}^{(t)}=\mathbb{I}(p_{i}^{(t)}\leq v_{i}^{(t)})$ is equal to $1$. Let $\mu_{i}=1-F_{i}$ and note that
$\mathbb{E}[y_{i}^{(t)}|\mathbf{p}^{(t)}]=\mu_{i}(p_{i}^{(t)}-\langle\widetilde{\boldsymbol{\theta}}_{i,0},\mathbf{p}^{(t)}\rangle)=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}^{(t)}\rangle),$
where $\boldsymbol{\theta}_{i,0}$ coincides with $\widetilde{\boldsymbol{\theta}}_{i,0}$ except for the $i$-th coordinate which instead is equal to $1-[\widetilde{\boldsymbol{\theta}}_{i,0}]_{i}$, where $\widetilde{\boldsymbol{\theta}}_{i,0}]_{i}$ is the $i$-th coordinate of $\widetilde{\boldsymbol{\theta}}_{i,0}$. This model coincides with the canonical exponential family in (2) with $\mu_{i}=b_{i}^{\prime}$.
Example 2.3 (Linear Regression Model).
Here we describe the model by Li et al. (2024). The observed demand at time $t$ for seller $i$ is assumed to have the form
$y_{i}^{(t)}=\alpha_{i,0}-\beta_{i,0}p_{i}^{(t)}+\langle\boldsymbol{\gamma}_{i,0},\mathbf{p}_{-i}^{(t)}\rangle+\eta_{i}^{(t)},$
$\boldsymbol{\eta}^{(t)}\mid\mathbf{p}^{(t)}\sim\mathcal{N}\left(\boldsymbol{0},\Sigma\left(\mathbf{p}^{(t)}\right)\right)$, with $\Sigma\left(\mathbf{p}^{(t)}\right)$ positive semidefinite, where $\boldsymbol{\eta}^{(t)}=(\eta_{i}^{(t)})_{i\in[N]}$. This model matches the canonical exponential family in (2) with $\mu_{i}(x)=b_{i}^{\prime}(x)=\alpha_{i,0}+x$.
We now present the assumptions on $\mu_{i}$ and the parameters $\boldsymbol{\theta}_{i,0}$. We assume that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\boldsymbol{\theta}_{i,0}\}_{i\in[N]}$; a similar assumption is found in Birge et al. (2024); Li et al. (2024); Bracale et al. (2025a). The above conditions hold if $\mu_{i},\mu_{i}^{\prime}>0$ and $\beta_{i,0}>0$, which are explicitly stated in 2.4 and subsequent 2.5.
Assumption 2.4 (Bounded and feasible parameter space).
The unknown vector $\boldsymbol{\theta}_{i,0}=(-\beta_{i,0},\boldsymbol{\gamma}_{i,0})\in\Theta_{i}\subset\mathbb{R}^{N}$ satisfies $0<\underline{\beta}_{i}\leq\beta_{i,0}\leq\bar{\beta}_{i}$, $\boldsymbol{\gamma}_{i,0}\in K_{i}\subset\mathbb{R}^{N-1}$ for some $K_{i}$ compact convex, so that $\Theta_{i}\triangleq[-\bar{\beta}_{i},-\underline{\beta}_{i}]\times K_{i}$ is compact convex in $\mathbb{R}^{N}$. $\Theta_{i}$ is known to seller $i$.
We denote by $B_{\theta_{i}}>0$ the $\ell^{2}$ upper bound on $\Theta_{i}$, that is $\lVert\boldsymbol{\theta}_{i}\rVert_{2}\leq B_{\theta_{i}}$ for all $\boldsymbol{\theta}_{i}\in\Theta_{i}$.
As far as $\mu_{i}=b_{i}^{\prime}$, we already know that it is non-decreasing because $b_{i}$ is convex. However, to derive concentration bounds, we require $\mu_{i}^{\prime}$ to be bounded away from zero. Before making this explicit, we need to define
$\mathcal{U}_{i}\triangleq[u^{\min}_{i},u^{\max}_{i}],$
where $u_{i}^{\min}\triangleq\inf_{(\boldsymbol{\theta}_{i},\mathbf{p})\in\Theta_{i}\times\mathcal{P}}\boldsymbol{\theta}_{i}^{\top}\mathbf{p}$, and $u_{i}^{\max}\triangleq\sup_{(\boldsymbol{\theta}_{i},\mathbf{p})\in\Theta_{i}\times\mathcal{P}}\boldsymbol{\theta}_{i}^{\top}\mathbf{p}$ (known to seller $i$ by 2.4).
Assumption 2.5 (Smoothness of $\mu_{i}$).
We assume that $\mu_{i}\in\mathrm{C}^{2}(\mathcal{U}_{i})$ with
$0\leq\mu_{i}(u)\leq B_{i},\quad\mu_{i}^{\prime}(u)\geq c_{\mu_{i}},\quad\forall u\in\mathcal{U}_{i},$
and $\mu_{i}^{\prime}(u)\leq L_{\mu_{i}}$, $\forall u\in\mathbb{R}$, for some $B_{i},c_{\mu_{i}},L_{\mu_{i}}>0$. The function $\mu_{i}$ is known to seller $i$ but unknown to competitors $j\in[N]\setminus\{i\}$.
Requiring $\mu_{i}^{\prime}$ to be bounded away from zero is standard: it prevents degeneracy in the curvature of the log-likelihood (equivalently, keeps the Fisher information well-conditioned) and underpins the concentration inequalities used in our analysis; similar assumptions can be found in Russac et al. (2019, 2020); Balabdaoui et al. (2019); Groeneboom and Hendrickx (2018). Let $B_{i}^{\prime\prime}>0$ be such that $|\mu_{i}^{\prime\prime}(u)|\leq B_{i}^{\prime\prime}$ for all $u\in\mathcal{U}_{i}$, which exists because $\mu_{i}\in\mathrm{C}^{2}(\mathcal{U}_{i})$.
We define the noise term as
$\eta_{i}^{(t)}\triangleq y_{i}^{(t)}-\mu_{i}(\langle\mathbf{p}^{(t)},\boldsymbol{\theta}_{i,0}\rangle),$
which satisfies the following Lemma 2.6, used to derive concentration bounds in our analysis.
Lemma 2.6.
Under 2.5, for every $i\in[N]$, $\eta_{i}^{(t)}$ is $L_{\mu_{i}}$-subgaussian conditionally on $\mathcal{H}_{i}^{(t-1)}$.
Remark 2.7 (Impossibility to learn competitors’ models).
Before proceeding to the next section, we highlight a key distinction between our framework and the MARL models discussed in Section 1.2. In most MARL settings, the realized rewards of all agents – in our setting, the realized demands of competing sellers – are observed by every participant. By contrast, our model assumes that these realized demands remain private. This privacy has an important implication: no seller can infer the policy of their competitors. To see this, recall that the demand model for seller $j$ is $\mathbb{E}[y_{j}^{(t)}|\mathbf{p}^{(t)}]=\mu_{j}(\langle\boldsymbol{\theta}_{j,0},\mathbf{p}^{(t)}\rangle)$ where $\boldsymbol{\theta}_{j,0}$ and $\mu_{j}$ are the parameter vector and link function of seller $j$, respectively, both unknown to competitors. At each time $t$, a competitor seller $i\neq j$ only observes $\mathcal{H}_{i}^{(t)}=\cup_{s\leq t}\{(y_{i}^{(s)},\mathbf{p}^{(s)})\}$ but does not have access to the demand realizations $\{y_{j}^{(s)}\}_{s\leq t}$ of seller $j$. Thus, seller $i$ has no information from which to infer their competitor parameters $(\boldsymbol{\theta}_{j,0},\mu_{j})$, and hence predict their pricing policy.
2.1 Regret and Nash equilibrium
Each seller competes with a dynamic optimal sequence of prices in hindsight while assuming that the other sellers would not have responded differently if this sequence of prices had been offered. Under such a dynamic benchmark, the objective of each seller $i\in[N]$ is to minimize the following regret metric in hindsight:
$\mathrm{Reg}_{i}(T)\triangleq\mathbb{E}\sum_{t=1}^{T}\operatorname{rev}_{i}(\Gamma_{i}(\mathbf{p}_{-i}^{(t)})|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p_{i}^{(t)}|\mathbf{p}_{-i}^{(t)})$
where for every $(p_{i},\mathbf{p}_{-i})\in\mathcal{P}$ we define
$\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})$ $\triangleq p_{i}\lambda_{i}(\mathbf{p})=p_{i}\mu_{i}(\langle\mathbf{p},\boldsymbol{\theta}_{i,0}\rangle),$ (3)
and for every $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$
$\Gamma_{i}(\mathbf{p}_{-i})\triangleq\arg\max_{p_{i}\in\mathcal{P}_{i}}\,\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i}),$ (4)
$\Gamma_{i}:\mathcal{P}_{-i}\rightarrow\mathcal{P}_{i}$ being the $i$-th seller’s best-response map.
Assumption 2.8 (Strong Concavity).
For every $i\in[N]$, $\operatorname{rev}_{i}(\cdot|\mathbf{p}_{-i})$ is strongly concave in $\mathcal{P}_{i}$, uniformly in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. [^2] There exists $\xi_{i}>0$ independent of $\mathbf{p}_{-i}$ such that for all $x,y\in\mathcal{P}_{i}$ and all $\mathbf{p}_{-i}$, $\operatorname{rev}_{i}(y|\mathbf{p}_{-i})-\operatorname{rev}_{i}(x|\mathbf{p}_{-i})\leq\big\langle\nabla_{x}\operatorname{rev}_{i}(x|\mathbf{p}_{-i}),y-x\big\rangle-\frac{\xi_{i}}{2}\|y-x\|_{2}^{2}$.
2.8 is common in the literature (Bracale et al., 2025a; Scutari et al., 2014; Li et al., 2024; Tsekrekos and Yannacopoulos, 2024) and includes linear demand models (Guda and Subramanian, 2026; Li et al., 2024), concave demand specifications, and $s_{i}$-concave demand functions with $s_{i}\geq 1$ (see Bracale et al. (2025a), Appendix I).
2.8 guarantees that the best-response map $\Gamma_{i}(\cdot)$ in (4) is well defined, i.e., admits a unique value for every $\mathbf{p}_{-i}$, which can be found by solving the first-order condition $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})=0$, w.r.t. $p_{i}$ and projecting the solution on $\mathcal{P}_{i}$. Additionally, 2.8 guarantees that the learning dynamics converge with sublinear sellers’ regrets. Indeed, as in Li et al. (2024); Bracale et al. (2025a), we control the individual regret upper bound by controlling the $L^{2}$ difference of the joint policies with respect to the NE, i.e. $\operatorname{Reg}_{i}(T)\lesssim\sum_{t=1}^{T}\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}$ where $\mathbf{p}^{*}$ is a NE (see Equation 15 for a detailed bound). A Nash equilibrium $\mathbf{p}^{*}=\left(p_{i}^{*}\right)_{i\in[N]}\in\mathcal{P}$ is defined as a price vector under which unilateral deviation is not profitable for any seller. Specifically,
$\operatorname{rev}_{i}(p_{i}^{*}|\mathbf{p}_{-i}^{*})\geq\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i}^{*}),\quad\forall p_{i}\in\mathcal{P}_{i},\quad\forall i\in[N],$
or, equivalently, $\mathbf{p}^{*}$ is a solution to the following fixed point equation:
$\mathbf{p}^{*}=\boldsymbol{\Gamma}(\mathbf{p}^{*})=(\Gamma_{1}(\mathbf{p}^{*}_{-1}),\dots,\Gamma_{N}(\mathbf{p}^{*}_{-N})),$ (5)
where $\boldsymbol{\Gamma}:\mathcal{P}\rightarrow\mathcal{P}$ is called best-response operator. The next result is an immediate consequence of Theorem 3 in Scutari et al. (2014).
Lemma 2.9 (Existence of NE).
Under assumptions (2.1), (2.4), (2.5) and (2.8), there exists a $\mathbf{p}^{*}\in\mathcal{P}$ satisfying the fixed point equation in (5).
Our next assumption guarantees the uniqueness of the NE.
Assumption 2.10 (Sufficiency for unique NE).
$\mu_{i}$ is log-concave in $\mathcal{U}_{i}$, and $\sup_{i\in[N]}\frac{\|\boldsymbol{\gamma}_{i,0}\|_{1}}{\beta_{i,0}}<1$.
The log-concavity assumption covers the linear demand model by Li et al. (2024) for which $\mu_{i}(u)=\alpha_{i,0}+u$ is log-concave. This assumption can be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in Fan et al. (2024), Assumption 1 in Chen and Farias (2018), Equation 1 in Cole and Roughgarden (2014) and Golrezaei et al. (2019); Javanmard (2017); Javanmard and Nazerzadeh (2019) which assume $\mu_{i}$ and $1-\mu_{i}$ to be log-concave, where $\mu_{i}$ is a survival function. The assumption in the model parameter constraint coincides with Assumption 2 in Li et al. (2024), and similar assumptions can be found in Bracale et al. (2025a); Kachani et al. (2007).
Lemma 2.11 (Contraction of the best-response operator).
If assumption (2.10) holds in addition to the assumptions in Lemma 2.9, then $\boldsymbol{\Gamma}$, defined in (5), is a contraction under the supremum norm with contraction constant $L_{\boldsymbol{\Gamma}}:=\sup_{i\in[N]}\nicefrac{{\|\boldsymbol{\gamma}_{i,0}\|_{1}}}{{\beta_{i,0}}}<1$.
Lemma 2.11 was proved by Bracale et al. (2025a) in the more general case of $s_{i}$-concave demand model for $s_{i}>-1$ (log-concavity is covered for $s_{i}=0$) and implies that the NE is unique by the Banach Contraction theorem. As in Li et al. (2024); Bracale et al. (2025b) the uniqueness of the NE entails sublinear individual regret for all sellers, as we will see in the proof of our main Theorem 4.1.
3 Algorithm
We adopt a similar algorithm to Russac et al. (2020) based on UCB of the revenue functions $\operatorname{rev}_{i}(\cdot|\mathbf{p}_{-i})$ for every $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. This method relies on a penalized log-likelihood approach that has a regularizing effect and avoids the need for common initial exploration across the sellers.
3.1 Penalized MLE
By the exponential model in (2), the (penalized) maximum likelihood estimator $\widehat{\boldsymbol{\theta}}_{i}^{(t)}$ based on the observed $y_{i}^{(1)},\dots,y_{i}^{(t-1)}$ and the selected prices $\mathbf{p}^{(1)},\dots,\mathbf{p}^{(t-1)}$ is defined as the maximizer of
$\sum_{s=1}^{t-1}\ln(\mathbb{P}_{\boldsymbol{\theta}_{i}}(y_{i}^{(s)}|\mathbf{p}^{(s)}))-\frac{\lambda}{2}\lVert\boldsymbol{\theta}_{i}\rVert_{2}^{2}$ (6)
$=\sum_{s=1}^{t-1}y_{i}^{(s)}\langle\boldsymbol{\theta}_{i},\mathbf{p}^{(s)}\rangle-b_{i}(\langle\boldsymbol{\theta}_{i},\mathbf{p}^{(s)}\rangle)+c_{i}(y_{i}^{(s)})-\frac{\lambda}{2}\lVert\boldsymbol{\theta}_{i}\rVert_{2}^{2}$
which is strictly concave because, by convexity of $b_{i}$, $\sum_{s=1}^{t-1}\ln(\mathbb{P}_{\boldsymbol{\theta}_{i}}(y_{i}^{(s)}|\mathbf{p}^{(s)}))$ is concave in $\boldsymbol{\theta}_{i}$. Define $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ as the unique solution of
$\sum_{s=1}^{t-1}\left\{y_{i}^{(s)}-\mu_{i}(\langle\mathbf{p}^{(s)},\boldsymbol{\theta}_{i}\rangle)\right\}\mathbf{p}^{(s)}-\lambda\boldsymbol{\theta}_{i}=0,$ (7)
where the LHS is the derivative of the penalized log-likelihood in (6). Following the notation in Russac et al. (2020), we introduce
$V_{i}^{(t-1)}\triangleq\sum_{s=1}^{t-1}\mathbf{p}^{(s)}\mathbf{p}^{(s)\top}+\frac{\lambda}{c_{\mu_{i}}}I_{N},$ (8)
and
$g_{i}^{(t-1)}(\boldsymbol{\theta}_{i})\triangleq\sum_{s=1}^{t-1}\mu_{i}(\langle\mathbf{p}^{(s)},\boldsymbol{\theta}_{i}\rangle)\mathbf{p}^{(s)}+\lambda\boldsymbol{\theta}_{i}.$
For fixed $i$ and $t$, $g_{i}^{(t-1)}$ denotes a surrogate function introduced to obtain an upper bound on the matrix $V_{i}^{(t-1)}$ – in the sense of partial ordering on the non-negative matrices. This bound plays a key role in controlling the upper confidence bound that will be formally defined in Section 3.2. At every time $t$, the $i$-th seller estimator is $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$, defined by
$\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}=\underset{\lVert\boldsymbol{\theta}_{i}\rVert_{2}\leq B_{\theta_{i}}}{\operatorname{argmin}}\lVert g_{i}^{(t-1)}(\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML})-g_{i}^{(t-1)}(\boldsymbol{\theta}_{i})\rVert_{V_{i}^{(t-1),-1}},$ (9)
where $V_{i}^{(t-1),-1}$ is the inverse of $V_{i}^{(t-1)}$. We need to consider $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ because $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ is not guaranteed to satisfy $\lVert\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}\rVert_{2}\leq B_{\theta_{i}}$. Here, $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ should be understood as a “projection” onto the admissible parameter set.
3.2 Pricing Policy via UCB
In this section, we describe how each seller $i$ determines their prices $p_{i}^{(t)}$ at a given time $t$. Fix $t\in[T]$ and let $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}=(-\widetilde{\beta}_{i}^{(t)\tt PML},\widetilde{\boldsymbol{\gamma}}_{i}^{(t)\tt PML})$ be the parameter estimate defined in (9). For any $(p_{i},\mathbf{p}_{-i})\in\mathcal{P}$, we define the estimated revenue as the plug-in estimator:
$\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i})$ $=p_{i}\mu_{i}(-\widetilde{\beta}_{i}^{(t)\tt PML}p_{i}+\langle\widetilde{\boldsymbol{\gamma}}_{i}^{(t)\tt PML},\mathbf{p}_{-i}\rangle)$
$=p_{i}\mu_{i}(\langle\mathbf{p},\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}\rangle).$
Following standard UCB-based approaches (Russac et al., 2020, 2019), we define the upper confidence bound (UCB) of the estimated revenue as
$\mathrm{UCB}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i}^{(t-1)})=\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i}^{(t-1)})+\sigma_{i}^{(t-1)}(p_{i},\delta),$ (10)
where for all $s=0,1,\dots,T-1$ we define
$\sigma_{i}^{(s)}(p_{i},\delta)\triangleq\rho_{i}^{(s)}(\delta)p_{i}\lVert(p_{i},\mathbf{p}_{-i}^{(s)})\rVert_{V_{i}^{(s),-1}},$
$\rho_{i}^{(s)}(\delta)\triangleq\frac{2L_{\mu_{i}}}{c_{\mu_{i}}}\bigg(\mathrm{c}^{(s)}_{i}(\delta)+B_{\theta_{i}}\sqrt{c_{\mu_{i}}\lambda}\bigg),$
$\mathrm{c}^{(s)}_{i}(\delta)\triangleq L_{\mu_{i}}\sqrt{2\log\left(\frac{T}{\delta}\right)+N\log\left(1+\frac{B_{p}^{2}s}{N\lambda}\right)}.$ (11)
We finally define $p_{i}^{(t)}=\widehat{\Gamma}_{i}^{(t)}(\mathbf{p}^{(t-1)}_{-i})$, where
$\widehat{\Gamma}_{i}^{(t)}(\mathbf{p}^{(t-1)}_{-i})\in\operatorname{argmax}_{p_{i}\in\mathcal{P}_{i}}\,\mathrm{UCB}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i}^{(t-1)}),$ (12)
which represents an optimistic approximation of the best-response map $\Gamma_{i}$ at $\mathbf{p}^{(t-1)}_{-i}$.
[…]
Input: Probability $\delta$, regularization $\lambda$, upper bounds $B_{p}$, $B_{\theta_{i}}$, $B_{i}$, $c_{\mu_{i}}$, and $L_{\mu_{i}}$.
Initialize: $\widehat{\boldsymbol{\theta}}_{i}^{(0)\tt PML}\in\Theta_{i}$, each seller starts with any $p_{i}^{(0)}\in\mathcal{P}_{i}$ and observes the corresponding $y_{i}^{(0)}$. At the end of time $t=0$ each seller $i$ observes $\mathbf{p}^{(0)}$ and sets $V_{i}^{(0)}=\mathbf{p}^{(0)}\mathbf{p}^{(0)\top}+\nicefrac{{\lambda}}{{c_{\mu_{i}}}}I_{N}$.
for $t=1$ to $T$ do
  $\bullet$ Each seller $i$ computes $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ according to (7). If $\lVert\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}\rVert_{2}\leq B_{\theta_{i}}$ let $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}=\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$. Else compute $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ defined in (9).
  $\bullet$ Each seller $i$ set $p_{i}^{(t)}\leftarrow\widehat{\Gamma}_{i}^{(t)}(\mathbf{p}^{(t-1)}_{-i})$, defined in (12).
  $\bullet$ Each seller $i$ observes $y_{i}^{(t)}$.
  $\bullet$ Each seller $i$ observes $\mathbf{p}^{(t)}$.
  $\bullet$ Each seller $i$ updates $V_{i}^{(t)}\leftarrow V_{i}^{(t-1)}+\mathbf{p}^{(t)}\mathbf{p}^{(t)\top}$.
end for
4 Regret Analysis
In this section, we present our main result, Theorem 4.1, which shows the $\widetilde{O}(\sqrt{T})$ bound on the individual regret.
[…]
Suppose assumptions (2.1), (2.4), (2.5), (2.8) and (2.10) hold. For every $i\in[N]$ we have that, if $\delta=\frac{1}{T^{\gamma}}$ for some fixed $\gamma>1$, then, as long as $\lambda$ varies between $O(1/\sqrt{T})$ and $O(1)$,
$\operatorname{Reg}_{i}(T)\leq C+O(N^{2}\sqrt{T}\log(T)),$
for some $C>0$ independent of $T$. An upper bound expression for $\operatorname{Reg}_{i}(T)$ depending on $\lambda,N$ and $T$ can be found in (30). Additionally we have
$\sum_{t=1}^{T}\mathbb{E}[\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}]=O(N^{2}\sqrt{T}\log(T)),$ (13)
where $\mathbf{p}^{(t)}=\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})=(\widehat{\Gamma}_{i}^{(t)}(\mathbf{p}_{-i}^{(t-1)}))_{i\in[N]}$, i.e. $p_{i}^{(t)}=\widehat{\Gamma}_{i}^{(t)}(\mathbf{p}_{-i}^{(t-1)})$, $\forall i\in[N]$.
Proof Sketch.
Using Taylor expansion around $\Gamma_{i}(\mathbf{p}_{-i}^{(t)})$, we can write
$\operatorname{rev}_{i}\big(\Gamma_{i}(\mathbf{p}_{-i}^{(t)})|\mathbf{p}_{-i}^{(t)}\big)-\operatorname{rev}_{i}\big(p_{i}^{(t)}|\mathbf{p}_{-i}^{(t)}\big)$
$\lesssim(\Gamma_{i}(\mathbf{p}_{-i}^{(t)})-p_{i}^{(t)})^{2}\leq\|\boldsymbol{\Gamma}(\mathbf{p}^{(t)})-\mathbf{p}^{(t)}\|^{2}_{2},$
and summing over $t$ and taking $\mathbb{E}$ both sides we have
$\textstyle\mathrm{Reg}_{i}(T)\lesssim\sum_{t=1}^{T}\mathbb{E}\|\boldsymbol{\Gamma}(\mathbf{p}^{(t)})-\mathbf{p}^{(t)}\|_{2}^{2}.$ (14)
Adding and subtracting the fixed point $\mathbf{p}^{*}$ from the RHS of (14) and using that $\mathbf{p}^{*}=\boldsymbol{\Gamma}(\mathbf{p}^{*})$ along with Lipschitzianity of $\boldsymbol{\Gamma}$ we get that $\mathrm{Reg}_{i}(T)$ is bounded by
$\textstyle\sum_{t=1}^{T}\mathbb{E}\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}=\sum_{t=1}^{T}\mathbb{E}\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\mathbf{p}^{*}\|_{2}^{2}.$ (15)
Now let $\sigma^{(t)}(\delta)\triangleq\sum_{i\in[N]}\sigma_{i}^{(t)}(p_{i}^{(t+1)},\delta)$. Adding and subtracting $\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})$ from $\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\mathbf{p}^{*}\|_{2}$ and iterating this inequality $T$ times, we get
$\textstyle\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\mathbf{p}^{*}\|\lesssim\left(\sum_{j=0}^{t-1}L^{j}_{\boldsymbol{\Gamma}}\sqrt{\sigma^{(t-j-1)}(\delta)}\right)+L^{t-1}_{\boldsymbol{\Gamma}},$ (16)
which holds with high probability by the following Proposition 4.2.
Proposition 4.2.
Let the assumptions of the theorem hold. For every $0<\delta<1$, the event
$\cap_{t\in[T]}\{\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})\|_{2}^{2}\lesssim\sigma^{(t-1)}(\delta)\},$
holds with probability at least $1-2N\delta$.
Plugging the bound from (16) into the expectation (15) (and ignoring the expectation over the complementary probability set, which can be shown to be negligible), and using that $L_{\boldsymbol{\Gamma}}<1$, we obtain
$\mathrm{Reg}_{i}(T)\lesssim\mathbb{E}\sum_{t=0}^{T-1}\sigma^{(t)}(\delta).$
Now, for simplicity of the proof’s sketch, consider $\lambda$ a fixed constant and choose $\delta=1/T^{2}$. From the definition of $\rho_{i}^{(t)}(\delta)$ in Section 3.2 it is easy to see that $\rho_{i}^{(t)}(\delta)\lesssim\sqrt{N\log(T)}$, hence
$\sigma^{(t)}(\delta)$ $=\sum_{j\in[N]}\rho_{j}^{(t)}(\delta)p_{j}^{(t+1)}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}$
$\lesssim\sqrt{N\log(T)}\sum_{j\in[N]}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}.$
Therefore we obtain
$\mathrm{Reg}_{i}(T)\lesssim\sqrt{N\log(T)}\cdot K(T),$ (17)
where $K(T)=\mathbb{E}\sum_{t=0}^{T-1}\sum_{j\in[N]}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}$. The reader can immediately recognize the correspondence of $K(T)$ with the elliptical potential lemma used in the bandit literature, with the only difference being that, in a multi-agent setting, the presence of $p_{j}^{(t+1)}$ does not allow a direct application of this lemma. However, in Lemma C.2 we prove that it is still possible to bound $K(T)$ by $O(N^{3/2}\sqrt{T\log(T)})$ without additional assumptions. This completes the regret convergence since, by Equation 17, we obtain $\mathrm{Reg}_{i}(T)\lesssim\sqrt{N\log(T)}\cdot K(T)=O(N^{2}\sqrt{T}\log(T))$. The convergence of $\sum_{t=1}^{T}\mathbb{E}\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}$ comes from (15). ∎
[…]
Remark 4.4 (Computational complexity.).
At round $t$, seller $i$ updates the matrix $V_{i}^{(t)}=V_{i}^{(t-1)}+\mathbf{p}^{(t)}\mathbf{p}^{(t)\top}$ and its inverse. Using the Sherman–Morrison formula to update $V_{i}^{(t),-1}$ costs $O(N^{2})$ time (and forming $V_{i}^{(t)}$ itself is also $O(N^{2})$). To evaluate the UCB bonus for a candidate price $p_{i}$, one needs $\|(p_{i},\mathbf{p}_{-i})\|_{V_{i}^{(t),-1}}$, which requires a matrix–vector product $V_{i}^{(t),-1}\mathbf{p}$ in $O(N^{2})$ time and then a dot product in $O(N)$. Both the true best response and the optimistic best response solve a one-dimensional concave maximization problem. In practice, this can be implemented via a 1-D convex optimization (e.g., bisection or Newton), which is costless, i.e., $O(1)$ (this implies that the UCB optimization is tractable and does not suffer from the NP-hardness of high-dimensional problems). Over $T$ rounds, the time is $O(TN^{2})$ per seller, and $O(TN^{3})$ if all $N$ sellers are run. Space usage per seller is $O(N^{2})$ to store $V_{i}^{(t)}$ and (optionally) $V_{i}^{(t),-1}$ plus $O(N)$ for current vectors; if past histories are not stored, $O(N^{2})$ dominates.
4.1 Convergence to Nash Equilibrium
The bound in (13) alone does not imply a convergence rate for $\mathbb{E}\left[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}\right]$ as $T\to\infty$. Indeed, a bound on the partial sums only yields pointwise convergence if the sequence is sufficiently regular (e.g., eventually monotone), or under an initial i.i.d. pricing exploration phase, like in Li et al. (2024) and Bracale et al. (2025a), which ensures that the estimation error satisfies $\|\widehat{\boldsymbol{\theta}}_{i}^{(T)}-\boldsymbol{\theta}_{i,0}\|_{2}=O_{P}(T^{-1/2})$ under standard regularity conditions, which in turn yields the parametric-rate convergence $\mathbb{E}\left[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}\right]=O(T^{-1/2})$.
In our setting, however, prices $\{\mathbf{p}^{(t)}\}_{t\leq T}$ are generated adaptively (their selection depends on past noise realizations), and we do not impose an i.i.d. excitation phase. In the absence of this phase, the estimation error can typically be controlled only along directions that are sufficiently excited by the played actions. As a result, $\|\widehat{\boldsymbol{\theta}}_{i}^{(T)}-\boldsymbol{\theta}_{i,0}\|_{2}$ does not generally decay at the parametric rate $O_{P}(T^{-1/2})$, and one should not expect a pointwise convergence rate for $\mathbb{E}\left[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}\right]$ based solely on (13).
Nevertheless, as we show in Theorem 4.1, our primary objective is regret optimality, which can be achieved even without strong pointwise convergence of the price sequence to a Nash equilibrium. This is consistent with the broader bandit literature, where regret-optimal algorithms often rely on adaptive (non-i.i.d.) action selection, as in OFU/UCB-style methods for linear and generalized linear bandits (Dani et al., 2008; Abbasi-Yadkori et al., 2011; Filippi et al., 2010; Russac et al., 2019).
5 Limitations and future directions.
We assume a known increasing link $\mu_{i}$, strong concavity of per-round revenues, and a contraction-type condition that guarantees a unique NE. None of these is new: they are standard in dynamic pricing and learning-in-games and strictly include the linear-demand setting of Li et al. (2024) as a special case. These conditions ensure well-posed best responses and well-conditioned information for concentration, but they also circumscribe our scope to stationary parameters and synchronous price moves under a unique-equilibrium regime.
A natural avenue is to relax link knowledge by replacing it with shape constraints (e.g., monotonicity or $s$-concavity) and to derive calibration-free confidence bonuses that adapt to local curvature while remaining free of any initial exploration phase. On the game-theoretic side, moving beyond global contractions – to allow multiple or set-valued equilibria – suggests a variational-inequality/monotone-operator treatment to obtain no-regret convergence. Algorithmically, Bayesian or Thompson sampling approaches to GLMs under our information structure (where rivals’ revenues remain unobserved) are particularly appealing, as they tend to yield less conservative confidence bounds in practice compared to UCB-based methods. Finally, incorporating nonstationarity in $\boldsymbol{\theta}_{i,0}$, inventory constraints, and asynchronous observability of rivals’ prices would bring the model even closer to more real market scenarios.
Appendix A Proof of Proposition 4.2
We want to prove that, under the assumptions of (4.1), for every $0<\delta<1$,
$\mathbb{P}\left(\cap_{t\in[T]}\left\{\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})\|_{2}^{2}\leq C_{2}\sigma^{(t-1)}(\delta)\right\}\right)\geq 1-2N\delta,$
where $\sigma^{(t)}(\delta)\triangleq\sum_{i\in[N]}\sigma_{i}^{(t)}(p_{i}^{(t+1)},\delta)$ and $C_{2}=\frac{4}{\min_{i\in[N]}\xi_{i}}$, with $\xi_{i}>0$ such that for all $x,y\in\mathcal{P}_{i}$ and all $\mathbf{p}_{-i}$,
$\operatorname{rev}_{i}(y|\mathbf{p}_{-i})-\operatorname{rev}_{i}(x|\mathbf{p}_{-i})\leq\big\langle\nabla_{x}\operatorname{rev}_{i}(x|\mathbf{p}_{-i}),y-x\big\rangle-\frac{\xi_{i}}{2}\|y-x\|^{2}$
For every $i\in[N]$ the constant $\xi_{i}$ is independent of $\mathbf{p}_{-i}$ by (2.8). We recall the definitions
$\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})=p_{i}\mu_{i}(-\beta_{i,0}p_{i}+\langle\boldsymbol{\gamma}_{i,0},\mathbf{p}_{-i}\rangle),\qquad\forall(p_{i},\mathbf{p}_{-i})\in\mathcal{P},$
and
$\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i})=p_{i}\mu_{i}(-\widetilde{\beta}_{i}^{(t)\tt PML}p_{i}+\langle\widetilde{\boldsymbol{\gamma}}_{i}^{(t)\tt PML},\mathbf{p}_{-i}\rangle),\qquad\forall(p_{i},\mathbf{p}_{-i})\in\mathcal{P}.$
For simplicity of notation, we also define
$p_{i}^{(t)\star}=\Gamma_{i}(\mathbf{p}^{(t)}_{-i})=\underset{p_{i}\in\mathcal{P}_{i}}{\operatorname{argmax}}\,\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i}^{(t)}),\qquad\forall\mathbf{p}_{-i}^{(t)}\in\mathcal{P}_{-i},$
and
$p_{i}^{(t+1)}=\widehat{\Gamma}^{(t+1)}_{i}(\mathbf{p}^{(t)}_{-i})\in\underset{p_{i}\in\mathcal{P}_{i}}{\operatorname{argmax}}\,\text{UCB}_{i}^{(t+1)}(\cdot|\mathbf{p}_{-i}^{(t)}),\qquad\forall\mathbf{p}_{-i}^{(t)}\in\mathcal{P}_{-i},$
where we recall from Equation 10 that
$\mathrm{UCB}_{i}^{(t+1)}(p_{i}|\mathbf{p}_{-i}^{(t)})=\widehat{\operatorname{rev}}_{i}^{(t+1)}(p_{i}|\mathbf{p}_{-i}^{(t)})+\sigma_{i}^{(t)}(p_{i},\delta).$
For the first part of the proof, we follow similar steps to Russac et al. (2020), Corollary 1. We start by upper-bounding the following difference:
$\operatorname{rev}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})$ $=\underbrace{\operatorname{rev}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})}_{A1}$
$+\underbrace{\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})}_{A2}$
$+\underbrace{\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})}_{A3}.$
To upper bound $A1$ and $A3$ we use the following Lemma A.1, whose proof can be found in Appendix B.
Lemma A.1.
Let assumptions (2.1), (2.4), (2.5) and (2.8) hold. Recall that for every $t=0,1,\dots,T-1$ we have defined
$\sigma_{i}^{(t)}(p_{i},\delta)=\rho_{i}^{(t)}(\delta)p_{i}\lVert(p_{i},\mathbf{p}_{-i}^{(t)})\rVert_{V_{i}^{(t),-1}},$
$\rho_{i}^{(t)}(\delta)=\frac{2L_{\mu_{i}}}{c_{\mu_{i}}}\bigg(\mathrm{c}^{(t)}_{i}(\delta)+B_{\theta_{i}}\sqrt{c_{\mu_{i}}\lambda}\bigg),\quad\text{and}\quad\mathrm{c}^{(t)}_{i}(\delta)=L_{\mu_{i}}\sqrt{2\log\left(\frac{T}{\delta}\right)+N\log\left(1+\frac{B_{p}^{2}t}{N\lambda}\right)}.$
Let $0<\delta<1$ and $t\in[T]$. Let $\mathbf{p}$ be any $\mathcal{P}$-valued (possibly random) price vector. Then, simultaneously for all $t\in[T]$,
$|\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})-\widehat{\operatorname{rev}}^{(t)}_{i}(p_{i}|\mathbf{p}_{-i})|=\big|p_{i}\mu_{i}(\langle\mathbf{p},\boldsymbol{\theta}_{i,0}\rangle)-p_{i}\mu_{i}(\langle\mathbf{p},\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}\rangle)\big|\leq\rho_{i}^{(t-1)}(\delta)p_{i}\lVert\mathbf{p}\rVert_{V_{i}^{(t-1),-1}}$
holds with probability higher than $1-\delta$.
Thanks to Lemma A.1, we can give an upper bound for the term $A1$ and for the term $A3$. With a union bound, we can simultaneously upper bound $A1$ and $A3$ for all $t\in[T]$ and the following holds
$\mathbb{P}\left(\forall t\in[T],A1\leq\rho_{i}^{(t)}(\delta)p_{i}^{(t)\star}\|(p_{i}^{(t)\star},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}\cap A3\leq\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\|(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}\right)\geq 1-2\delta$ (18)
Let $E$ be the event in (18). We now upper-bound $A2$.
$A2$ $=\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})$
$=\underset{=\mathrm{UCB}_{i}^{(t+1)}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})}{\underbrace{\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})+\rho_{i}^{(t)}(\delta)p_{i}^{(t)\star}\|(p_{i}^{(t)\star},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}}}$
$\qquad-\rho_{i}^{(t)}(\delta)p_{i}^{(t)\star}\|(p_{i}^{(t)\star},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}+\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\lVert(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\rVert_{V_{i}^{(t),-1}}$
$\qquad\underset{=-\mathrm{UCB}_{i}^{(t+1)}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})}{\underbrace{-\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\lVert(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\rVert_{V_{i}^{(t),-1}}-\widehat{\operatorname{rev}}^{(t+1)}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})}}$
$=\underset{\leq 0\text{ by def of }p_{i}^{(t+1)}}{\underbrace{\mathrm{UCB}_{i}^{(t+1)}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\mathrm{UCB}_{i}^{(t+1)}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})}}$
$\quad-\rho_{i}^{(t)}(\delta)p_{i}^{(t)\star}\|(p_{i}^{(t)\star},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}+\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\lVert(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\rVert_{V_{i}^{(t),-1}}$
$\leq\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\|(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}-\rho_{i}^{(t)}(\delta)p_{i}^{(t)\star}\|(p_{i}^{(t)\star},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}.$
Now we put $A1,A2,A3$ together. Under the event $E$ (which occurs with a probability higher than $1-2\delta$), we have
$\operatorname{rev}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})$ $\leq\underbrace{\rho_{i}^{(t)}(\delta)p_{i}^{(t)\star}\|(p_{i}^{(t)\star},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}}_{\text{coming from}A1}$
$\qquad+\underbrace{\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\|(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}-\rho_{i}^{(t)}(\delta)p_{i}^{(t)\star}\|(p_{i}^{(t)\star},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}}_{\text{coming from}A2}$
$\qquad+\underbrace{\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\|(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}}_{\text{coming from}A3}$
$\leq 2\rho_{i}^{(t)}(\delta)p_{i}^{(t+1)}\|(p_{i}^{(t+1)},\mathbf{p}_{-i}^{(t)})\|_{V_{i}^{(t),-1}}=2\sigma_{i}^{(t)}(p_{i}^{(t+1)},\delta).$
Up to here we proved that for $0<\delta<1$ and for every $i\in[N]$,
$\mathbb{P}\left(\cap_{t\in[T]}\left\{\operatorname{rev}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})\leq 2\sigma_{i}^{(t)}(p_{i}^{(t+1)},\delta)\right\}\right)\geq 1-2\delta.$ (19)
Now, by 2.8 there exists $\xi_{i}>0$ independent of $\mathbf{p}_{-i}^{(t)}$ such that for all $x,y\in\mathcal{P}_{i}$ and all $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$,
$\operatorname{rev}_{i}(y|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(x|\mathbf{p}_{-i}^{(t)})\leq\big\langle\nabla_{x}\operatorname{rev}_{i}(x|\mathbf{p}_{-i}^{(t)}),y-x\big\rangle-\frac{\xi_{i}}{2}\|y-x\|^{2}.$
replacing $x$ with $p_{i}^{(t)*}$ and $y$ with $p_{i}^{(t+1)}$, together with the optimality of $p_{i}^{(t)*}$, we recover that
$(p_{i}^{(t)*}-p_{i}^{(t+1)})^{2}\leq\frac{2}{\xi_{i}}[\operatorname{rev}_{i}(p_{i}^{(t)\star}|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p_{i}^{(t+1)}|\mathbf{p}_{-i}^{(t)})]$ (20)
Now define $C_{i,2}\triangleq\frac{4}{\xi_{i}}$, and the event
$\mathcal{B}_{i}\triangleq\cap_{t=1}^{T}\left\{(\Gamma_{i}(\mathbf{p}^{(t)}_{-i})-\widehat{\Gamma}_{i}^{(t+1)}(\mathbf{p}^{(t)}_{-i}))^{2}\leq C_{i,2}\sigma_{i}^{(t)}(p_{i}^{(t+1)},\delta)\right\}.$
Since $(\Gamma_{i}(\mathbf{p}^{(t)}_{-i})-\widehat{\Gamma}_{i}^{(t+1)}(\mathbf{p}^{(t)}_{-i}))^{2}=(p_{i}^{(t)*}-p_{i}^{(t+1)})^{2}$, then (19) and (20) imply that for $\delta\in(0,1)$ and for every $i\in[N]$, $\mathbb{P}(\mathcal{B}_{i})\geq 1-2\delta$. Now define $\mathcal{B}=\cap_{i\in[N]}\mathcal{B}_{i}$. Note that $\mathcal{B}$ holds with probability at least $1-2N\delta$:
$\mathbb{P}(\mathcal{B}^{c})=\mathbb{P}(\cup_{i\in[N]}\mathcal{B}_{i}^{c})\leq\sum_{i\in[N]}\mathbb{P}(\mathcal{B}_{i}^{c})=\sum_{i\in[N]}(1-\mathbb{P}(\mathcal{B}_{i}))\leq\sum_{i\in[N]}2\delta=2N\delta.$
Now define
$C_{2}\triangleq\max_{i\in[N]}\{C_{i,2}\}=\frac{4}{\min_{i\in[N]}\xi_{i}},\qquad\sigma^{(t)}(\delta)\triangleq\sum_{i\in[N]}\sigma_{i}^{(t)}(p_{i}^{(t+1)},\delta),\qquad\forall t\geq 1.$
In $\mathcal{B}$ we have that
$\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})\|_{2}$ $=\sqrt{\sum_{i\in[N]}(\widehat{\Gamma}^{(t)}_{i}(\mathbf{p}^{(t-1)}_{-i})-\Gamma_{i}(\mathbf{p}^{(t-1)}_{-i}))^{2}}$
$\leq\sqrt{\max_{i\in[N]}\{C_{i,2}\}}\sqrt{\sum_{i\in[N]}\sigma_{i}^{(t-1)}(\delta)}$
$=\sqrt{C_{2}\sigma^{(t-1)}(\delta)}.$
Let
$\mathcal{G}=\cap_{t\in[T]}\left\{\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})\|_{2}^{2}\leq C_{2}\sigma^{(t-1)}(\delta)\right\}.$
Since $\mathcal{B}$ implies $\mathcal{G}$, we have
$\mathbb{P}\left(\mathcal{G}\right)\geq\mathbb{P}\left(\mathcal{B}\right)\geq 1-2N\delta.$
This completes the proof.
Appendix B Proof of Lemma A.1
For simplicity of notation, we write $\widehat{\boldsymbol{\theta}}_{i}^{(t)}$ for $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ and $\widetilde{\boldsymbol{\theta}}_{i}^{(t)}$ for $\widetilde{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$. This proof follows a similar structure to the proof of Proposition 1 in Russac et al. (2020). Recall the definitions
$\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})=p_{i}\mu_{i}(-\beta_{i,0}p_{i}+\langle\boldsymbol{\gamma}_{i,0},\mathbf{p}_{-i}\rangle)=p_{i}\mu_{i}(\langle\mathbf{p},\boldsymbol{\theta}_{i,0}\rangle),$
and
$\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i})=p_{i}\mu_{i}(-\widetilde{\beta}_{i}^{(t)\tt PML}p_{i}+\langle\widetilde{\boldsymbol{\gamma}}_{i}^{(t)\tt PML},\mathbf{p}_{-i}\rangle)=p_{i}\mu_{i}(\langle\mathbf{p},\widetilde{\boldsymbol{\theta}}_{i}^{(t)}\rangle).$
We recall the definition $g_{i}^{(t-1)}:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$:
$g_{i}^{(t-1)}(\boldsymbol{\theta}_{i})=\sum_{s=1}^{t-1}\mu_{i}(\mathbf{p}^{(s)\top}\boldsymbol{\theta}_{i})\mathbf{p}^{(s)}+\lambda\boldsymbol{\theta}_{i}.$
Let $J_{i}^{(t-1)}$ denote the Jacobian matrix of $g_{i}^{(t-1)}$. We have
$J_{i}^{(t-1)}(\boldsymbol{\theta}_{i})=\sum_{s=1}^{t-1}\mu_{i}^{\prime}(\mathbf{p}^{(s)\top}\boldsymbol{\theta}_{i})\mathbf{p}^{(s)}\mathbf{p}^{(s)\top}+\lambda I_{N}.$
Thanks to the definition of the estimator $\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML}$ defined in Equation 7, we have
$g_{i}^{(t-1)}(\widehat{\boldsymbol{\theta}}_{i}^{(t)\tt PML})=\sum_{s=1}^{t-1}\mathbf{p}^{(s)}y_{i}^{(s)}.$
We also introduce the martingale
$S_{i}^{(t-1)}=\sum_{s=1}^{t-1}\mathbf{p}^{(s)}\eta_{i}^{(s)}.$
We define the $G_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0},\widetilde{\boldsymbol{\theta}}_{i}^{(t)})$ matrix as follows,
$G_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0},\widetilde{\boldsymbol{\theta}}_{i}^{(t)})=\int_{0}^{1}J_{i}^{(t-1)}(u\boldsymbol{\theta}_{i,0}+(1-u)\widetilde{\boldsymbol{\theta}}_{i}^{(t)})du.$
The Fundamental Theorem of Calculus gives
$g_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0})-g_{i}^{(t-1)}(\widetilde{\boldsymbol{\theta}}_{i}^{(t)})=G_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0},\widetilde{\boldsymbol{\theta}}_{i}^{(t)})(\boldsymbol{\theta}_{i,0}-\widetilde{\boldsymbol{\theta}}_{i}^{(t)}).$ (21)
Knowing that both $\boldsymbol{\theta}_{i,0}$ and $\widetilde{\boldsymbol{\theta}}_{i}^{(t)}$ have an L2-norm smaller than $B_{\theta_{i}}$, $\forall u\in[0,1],\lVert u\boldsymbol{\theta}_{i,0}+(1-u)\widetilde{\boldsymbol{\theta}}_{i}^{(t)})\rVert_{2}\leq B_{\theta_{i}}$. This implies in particular that
$G_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0},\widetilde{\boldsymbol{\theta}}_{i}^{(t)})\geq c_{\mu_{i}}\left(\sum_{s=1}^{t-1}\mathbf{p}^{(s)}\mathbf{p}^{(s)\top}+\frac{\lambda}{c_{\mu_{i}}}I_{N}\right)=c_{\mu_{i}}V_{i}^{(t-1)},$ (22)
which in turn ensures $G_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0},\widetilde{\boldsymbol{\theta}}_{i}^{(t)})$ is invertible. Let $\mathbf{p}$ be any price vector in $\mathcal{P}$ (possibly random) and $t$ be a fixed time instant,
$|\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})-\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i})|\leq p_{i}|\mu_{i}(\langle\mathbf{p},\boldsymbol{\theta}_{i,0}\rangle)-\mu_{i}(\langle\mathbf{p},\widetilde{\boldsymbol{\theta}}_{i}^{(t)}\rangle)|$ (23)
where
$|\mu_{i}(\langle\mathbf{p},\boldsymbol{\theta}_{i,0}\rangle)-\mu_{i}(\langle\mathbf{p},\widetilde{\boldsymbol{\theta}}_{i}^{(t)}\rangle)|$
[…]
$\leq L_{\mu_{i}}\lVert\mathbf{p}\rVert_{G_{i}^{(t-1),-1}(\boldsymbol{\theta}_{i,0},\widetilde{\boldsymbol{\theta}}_{i}^{(t)})}\lVert g_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0})-g_{i}^{(t-1)}(\widetilde{\boldsymbol{\theta}}_{i}^{(t)})\rVert_{G_{i}^{(t-1),-1}(\boldsymbol{\theta}_{i,0},\widetilde{\boldsymbol{\theta}}_{i}^{(t)})}\quad\text{(by Cauchy-Schwartz inequality)}$
[…]
$\leq\frac{2L_{\mu_{i}}}{c_{\mu_{i}}}\lVert\mathbf{p}\rVert_{V_{i}^{(t-1),-1}}\lVert g_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0})-g_{i}^{(t-1)}(\widehat{\boldsymbol{\theta}}_{i}^{(t)})\rVert_{V_{i}^{(t-1),-1}}\quad\text{(by definition of $\widetilde{\boldsymbol{\theta}}_{i}^{(t)}$)}.$
Now we calculate
$\lVert g_{i}^{(t-1)}(\boldsymbol{\theta}_{i,0})-g_{i}^{(t-1)}(\widehat{\boldsymbol{\theta}}_{i}^{(t)})\rVert_{V_{i}^{(t-1),-1}}$ $\leq\left\|\sum_{s=1}^{t-1}\mu_{i}(\mathbf{p}^{(s)\top}\boldsymbol{\theta}_{i,0})\mathbf{p}^{(s)}+\lambda\boldsymbol{\theta}_{i,0}-\sum_{s=1}^{t-1}\mathbf{p}^{(s)}y_{i}^{(s)}\right\|_{V_{i}^{(t-1),-1}}$
$\leq\left\|-\sum_{s=1}^{t-1}\mathbf{p}^{(s)}\eta_{i}^{(s)}+\lambda\boldsymbol{\theta}_{i,0}\right\|_{V_{i}^{(t-1),-1}}$
$=\left\|-S_{i}^{(t-1)}+\lambda\boldsymbol{\theta}_{i,0}\right\|_{V_{i}^{(t-1),-1}}$
$\leq\lVert S_{i}^{(t-1)}\rVert_{V_{i}^{(t-1),-1}}+\lVert\lambda\boldsymbol{\theta}_{i,0}\rVert_{V_{i}^{(t-1),-1}}\quad\text{(Triangle inequality)}$
$\leq\lVert S_{i}^{(t-1)}\rVert_{V_{i}^{(t-1),-1}}+\sqrt{\lambda c_{\mu_{i}}}\lVert\boldsymbol{\theta}_{i,0}\rVert_{2}\quad\left(V_{i}^{(t-1)}\geq\frac{\lambda}{c_{\mu_{i}}}I_{N}\right)$
$\leq\underset{=\mathrm{c}_{i}^{(t-1)}(\delta)}{\underbrace{L_{\mu_{i}}\sqrt{2\log(T/\delta)+N\log\left(1+\frac{B_{p}^{2}(t-1)}{N\lambda}\right)}}}+B_{\theta_{i}}\sqrt{\lambda c_{\mu_{i}}}\quad(\text{w.p.}\geq 1-\delta).$
In the last inequality, we have used the assumption that $\lVert\boldsymbol{\theta}_{i,0}\rVert_{2}\leq B_{\theta_{i}}$ and the concentration result established in Proposition 7 of Russac et al. (2019) for the self-normalized quantity $\|S_{i}^{(t-1)}\|_{V_{i}^{(t-1),-1}}$ (which can be applied thanks to the conditional $L_{\mu_{i}}$-subgaussianity of $\eta_{i}^{(t)}$ established in Lemma 2.6333In Russac et al. (2019), the subgaussianity of the error term is the sole assumption required to prove their Proposition 5, which in turn implies Proposition 7.), which establishes that
$\mathbb{P}\left(\exists t\leq T,\|S_{i}^{(t)}\|_{V^{(t),{-1}}}\geq\mathrm{c}^{(t)}_{i}(\delta)\right)\leq\delta,\quad\text{equivalently}\quad\mathbb{P}\left(\|S_{i}^{(t)}\|_{V^{(t),{-1}}}\leq\mathrm{c}^{(t)}_{i}(\delta),\forall t\leq T\right)\geq 1-\delta.$
We established that, for any $\mathbf{p}\in\mathcal{P}$,
$\mathbb{P}\left(|\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})-\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i})|\leq p_{i}\lVert\mathbf{p}\rVert_{V_{i}^{(t-1),-1}}\cdot\underset{=\rho_{i}^{(t-1)}(\delta)}{\underbrace{\frac{2L_{\mu_{i}}}{c_{\mu_{i}}}\left(\mathrm{c}_{i}^{(t-1)}(\delta)+B_{\theta_{i}}\sqrt{\lambda c_{\mu_{i}}}\right)}},\quad\forall t\leq T\right)\geq 1-\delta.$
This completes the proof.
Appendix C Proof of Theorem 4.1
We recall the definitions
$\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})=p_{i}\mu_{i}(-\beta_{i,0}p_{i}+\langle\boldsymbol{\gamma}_{i,0},\mathbf{p}_{-i}\rangle),\qquad\forall(p_{i},\mathbf{p}_{-i})\in\mathcal{P},$
and
$\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i})=p_{i}\mu_{i}(-\widetilde{\beta}_{i}^{(t)\tt PML}p_{i}+\langle\widetilde{\boldsymbol{\gamma}}_{i}^{(t)\tt PML},\mathbf{p}_{-i}\rangle),\qquad\forall(p_{i},\mathbf{p}_{-i})\in\mathcal{P}.$
For simplicity of notation, we also define
$p_{i}^{(t)\star}=\Gamma_{i}(\mathbf{p}^{(t)}_{-i})=\underset{p_{i}\in\mathcal{P}_{i}}{\operatorname{argmax}}\,\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i}^{(t)}),\qquad\forall\mathbf{p}_{-i}^{(t)}\in\mathcal{P}_{-i},$
and
$p_{i}^{(t)}=\widehat{\Gamma}^{(t)}_{i}(\mathbf{p}^{(t-1)}_{-i})\in\underset{p_{i}\in\mathcal{P}_{i}}{\operatorname{argmax}}\,\text{UCB}_{i}^{(t)}(\cdot|\mathbf{p}_{-i}^{(t-1)}),\qquad\forall\mathbf{p}_{-i}^{(t-1)}\in\mathcal{P}_{-i},$
where we recall from Equation 10 that
$\mathrm{UCB}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i}^{(t-1)})=\widehat{\operatorname{rev}}_{i}^{(t)}(p_{i}|\mathbf{p}_{-i}^{(t-1)})+\sigma_{i}^{(t-1)}(p_{i},\delta).$
The goal is to upper bound the individual regret
$\operatorname{Reg}_{i}(T)=\mathbb{E}\sum_{t=1}^{T}R_{i}^{(t)}=\mathbb{E}\sum_{t=1}^{T}[\operatorname{rev}_{i}(\Gamma_{i}(\mathbf{p}^{(t)}_{-i})|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p^{(t)}_{i}|\mathbf{p}_{-i}^{(t)})].$
We have
$R_{i}^{(t)}=\operatorname{rev}_{i}(\Gamma_{i}(\mathbf{p}^{(t)}_{-i})|\mathbf{p}_{-i}^{(t)})-\operatorname{rev}_{i}(p^{(t)}_{i}|\mathbf{p}_{-i}^{(t)})\leq C_{i,1}(\Gamma_{i}(\mathbf{p}^{(t)})-p_{i}^{(t)})^{2},$ (24)
where we used that $R_{i}^{(t)}\leq M_{i}$ and that
$|\partial^{2}_{p^{2}}\operatorname{rev}_{i}(p|\mathbf{p}_{-i}^{(t)})_{|p=p^{\prime}}|$ $=|-2\beta_{i,0}\mu_{i}^{\prime}(-\beta_{i,0}p^{\prime}+\boldsymbol{\gamma}_{i,0}^{\top}\mathbf{p}_{-i}^{(t)})+\beta_{i,0}^{2}p^{\prime}\mu_{i}^{\prime\prime}(-\beta_{i,0}p^{\prime}+\boldsymbol{\gamma}_{i,0}^{\top}\mathbf{p}_{-i}^{(t)})|$
$\leq 2\beta_{i,0}L_{\mu_{i}}+\beta_{i,0}^{2}\bar{p_{i}}B^{\prime\prime}_{i}$
$\leq 2\overline{\beta}_{i}L_{\mu_{i}}+\overline{\beta}_{i}^{2}\bar{p_{i}}B^{\prime\prime}_{i}\triangleq C_{i,1},$
for some $p^{\prime}$ in the segment between the points $p_{i}^{(t)}$ and $\Gamma_{i}(\mathbf{p}^{(t)}_{-i})$. Continuing from (24) we have
$(\Gamma_{i}(\mathbf{p}^{(t)}_{-i})-p_{i}^{(t)})^{2}$ $\leq\|\boldsymbol{\Gamma}(\mathbf{p}^{(t)})-\mathbf{p}^{(t)}\|_{2}^{2}$
$\leq 2\|\boldsymbol{\Gamma}(\mathbf{p}^{(t)})-\mathbf{p}^{*}\|_{2}^{2}+2\|\mathbf{p}^{*}-\mathbf{p}^{(t)}\|_{2}^{2}$
[…]
$\leq 2(L_{\boldsymbol{\Gamma}}+1)\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}.$ (25)
STEP 1: Bound for $\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}$.
Consider the event
$\mathcal{G}=\cap_{t\in[T]}\left\{\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})\|_{2}^{2}\leq C_{2}\sigma^{(t-1)}(\delta)\right\},$ (26)
that has probability at least $1-2N\delta$ by Proposition 4.2. Continuing from (C), we get that, on $\mathcal{G}$,
$\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}$ $=\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\mathbf{p}^{*}\|_{2}$
$\overset{(\star)}{\leq}\|\widehat{\boldsymbol{\Gamma}}^{(t)}(\mathbf{p}^{(t-1)})-\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})\|_{2}+\|\boldsymbol{\Gamma}(\mathbf{p}^{(t-1)})-\boldsymbol{\Gamma}(\mathbf{p}^{*})\|_{2}$
$\leq\sqrt{C_{2}\sigma^{(t-1)}(\delta)}+L_{\boldsymbol{\Gamma}}\|\mathbf{p}^{(t-1)}-\mathbf{p}^{*}\|_{2}$
$\leq\sqrt{C_{2}\sigma^{(t-1)}(\delta)}+L_{\boldsymbol{\Gamma}}\left(\sqrt{C_{2}\sigma^{(t-2)}(\delta)}+L_{\boldsymbol{\Gamma}}\|\mathbf{p}^{(t-2)}-\mathbf{p}^{*}\|_{2}\right)$
$=\sqrt{C_{2}}\left(\sqrt{\sigma^{(t-1)}(\delta)}+L_{\boldsymbol{\Gamma}}\sqrt{\sigma^{(t-2)}(\delta)}\right)+L^{2}_{\boldsymbol{\Gamma}}\|\mathbf{p}^{(t-2)}-\mathbf{p}^{*}\|_{2}$
$=\sqrt{C_{2}}\left(\sum_{j=0}^{\ell}L^{j}_{\boldsymbol{\Gamma}}\sqrt{\sigma^{(t-j-1)}(\delta)}\right)+L^{\ell}_{\boldsymbol{\Gamma}}\|\mathbf{p}^{(t-\ell)}-\mathbf{p}^{*}\|_{2},\qquad\ell\in\{1,\dots,t-1\}$
$\leq\sqrt{C_{2}}\left(\sum_{j=0}^{t-1}L^{j}_{\boldsymbol{\Gamma}}\sqrt{\sigma^{(t-j-1)}(\delta)}\right)+L^{t-1}_{\boldsymbol{\Gamma}}\|\mathbf{p}^{(0)}-\mathbf{p}^{*}\|_{2}.$ (27)
Summing over $t$ we obtain, on $\mathcal{G}$, the following inequality
$\sum_{t=1}^{T}\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}$ $\leq\sum_{t=1}^{T}\left\{2C_{2}\left(\sum_{j=0}^{t-1}L^{j}_{\boldsymbol{\Gamma}}\sqrt{\sigma^{(t-j-1)}(\delta)}\right)^{2}+2L^{2(t-1)}_{\boldsymbol{\Gamma}}\|\mathbf{p}^{(0)}-\mathbf{p}^{*}\|_{2}^{2}\right\}$
$\overset{(\star\star)}{\leq}\frac{2C_{2}}{(1-L_{\boldsymbol{\Gamma}})^{2}}\sum_{t=0}^{T-1}\sigma^{(t)}(\delta)+2\left(\sum_{t=1}^{T}L^{2(t-1)}_{\boldsymbol{\Gamma}}\right)\|\mathbf{p}^{(0)}-\mathbf{p}^{*}\|_{2}^{2}$
$\leq\frac{2C_{2}}{(1-L_{\boldsymbol{\Gamma}})^{2}}\sum_{t=0}^{T-1}\sigma^{(t)}(\delta)+\frac{2\|\mathbf{p}^{(0)}-\mathbf{p}^{*}\|_{2}^{2}}{1-L^{2}_{\boldsymbol{\Gamma}}}$
$=C_{4}\sum_{t=0}^{T-1}\sum_{j\in[N]}2\rho_{j}^{(t)}(\delta)p_{j}^{(t+1)}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}+C_{3},$ (28)
where $C_{3}=2\frac{\|\mathbf{p}^{(0)}-\mathbf{p}^{*}\|_{2}^{2}}{1-L^{2}_{\boldsymbol{\Gamma}}}$ and $C_{4}=\frac{2C_{2}}{(1-L_{\boldsymbol{\Gamma}})^{2}}$. In $(\star\star)$ we used the following Lemma C.1, by substituting $q\leftarrow L_{\Gamma}$ and $a_{k}\leftarrow\sqrt{\sigma^{(k)}(\delta)}$ (the proof is deferred to Appendix D).
Lemma C.1.
Let $(a_{k})_{k\geq 0}$ be any real sequence, $0<q<1$, and $T\geq 1$. Then
$\sum_{t=1}^{T}\Biggl(\sum_{j=0}^{t-1}q^{j}a_{t-j-1}\Biggr)^{2}\leq\frac{1}{(1-q)^{2}}\sum_{k=0}^{T-1}a_{k}^{2}.$
STEP 2: Regret Bound.
Consider the set $\mathcal{G}$ defined in (26) and the partitioning of the sampling space: $\mathcal{G}\cup\mathcal{G}^{c}$. We can write
$\mathbb{E}\left[R_{i}^{(t)}\right]$ $=\mathbb{E}\left[R_{i}^{(t)}\cdot\left(\mathbb{I}(\mathcal{G})+\mathbb{I}\left(\mathcal{G}^{c}\right)\right)\right]$
$\leq\mathbb{E}\left[R_{i}^{(t)}\cdot\mathbb{I}(\mathcal{G})\right]+M_{i}\mathbb{P}\left(\mathcal{G}^{c}\right)$
$\leq\mathbb{E}\left[R_{i}^{(t)}\cdot\mathbb{I}(\mathcal{G})\right]+M_{i}2N\delta.$
Where $M_{i}$ is an upper bound of $R_{i}^{(t)}$. For example, using that $0\leq\mu_{i}\leq B_{i}$ and $0\leq p_{i}\leq\bar{p}_{i}$, we can take $M_{i}=2B_{i}\bar{p}_{i}$. Then
$\mathbb{E}\left[\sum_{t=1}^{T}R_{i}^{(t)}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}R_{i}^{(t)}\cdot\mathbb{I}(\mathcal{G})\right]+M_{i}2N\delta T.$
Now, choosing $\delta=1/T^{\gamma}$ with $\gamma>1$, the second term converges to zero, so we only need to bound the first term on $\mathcal{G}$. Merging (24) and (C) we have that, on $\mathcal{G}$, up to a multiplicative constant $C_{4}$ and an additive constant $C_{3}$
$\operatorname{Reg}_{i}(T)=\mathbb{E}\left[\sum_{t=1}^{T}R_{i}^{(t)}\cdot\mathbb{I}(\mathcal{G})\right]$ $\lesssim\mathbb{E}\left[\sum_{t=1}^{T}\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}\cdot\mathbb{I}(\mathcal{G})\right]$
[…]
$\leq\bar{p}\,\mathbb{E}\left[\sum_{t=0}^{T-1}\sum_{j\in[N]}\rho_{j}^{(t)}(\delta)\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}\cdot\mathbb{I}(\mathcal{G})\right],$ (29)
where $\bar{p}=\max_{j\in[N]}\bar{p}_{j}>0$ and where we used that $\rho_{j}^{(t)}$ (defined in (3.2)) is increasing in $t$. Moreover, for $\delta=1/T^{\gamma}$ with $\gamma>1$, (retaining the dependence on $\lambda,N$ and $T$), we have
$\rho_{j}^{(t)}(\delta)=O\left(\mathrm{c}^{(t)}_{i}(\delta)+\sqrt{\lambda}\right)$ $=O\left(\sqrt{\lambda}+L_{\mu_{i}}\sqrt{2\log\left(\frac{T}{\delta}\right)+N\log\left(1+\frac{B_{p}^{2}T}{N\lambda}\right)}\right)$
$=O\left(\sqrt{\lambda}+\sqrt{\log(T)+N\log\left(1+\frac{T}{N\lambda}\right)}\right).$
It only remains to bound
$\sum_{t=0}^{T-1}\sum_{j\in[N]}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}},$
for which we use the following Lemma C.2, proved in Appendix E.
Lemma C.2 (Variant of the Elliptical Potential Lemma).
Retaining only the dependence on $\lambda,N$ and $T$, it holds
$\sum_{t=0}^{T-1}\sum_{j\in[N]}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}=O\left(\frac{N}{\sqrt{\lambda}}+N\sqrt{\lambda+T}+N^{3/2}\sqrt{T\log\left(\frac{T}{N\lambda}+1\right)}\right).$
We then conclude that the regret, in terms of $\lambda,N$ and $T$ is
$\operatorname{Reg}_{i}(T)=O\left(\left[\sqrt{\lambda}+\sqrt{\log(T)+N\log\left(1+\frac{T}{N\lambda}\right)}\right]\cdot\left[\frac{N}{\sqrt{\lambda}}+N\sqrt{\lambda+T}+N^{3/2}\sqrt{T\log\left(\frac{T}{N\lambda}+1\right)}\right]\right).$ (30)
As long as $\lambda$ varies between $O(1/\sqrt{T})$ and $O(1)$, then
$\operatorname{Reg}_{i}(T)=O\left(\sqrt{N\log(T)}\cdot N^{3/2}\sqrt{T\log(T)}\right)=O\left(N^{2}\sqrt{T}\log(T)\right).$
Note that from (C) we can also recover convergence to NE.
$\sum_{t=1}^{T}\mathbb{E}[\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}]=O(N^{2}\sqrt{T}\log(T)).$
This completes the proof.
Appendix D Proof of Lemma C.1
We want to prove that for a given sequence $(a_{k})_{k\geq 0}$ of real values, $0<q<1$, and $T\geq 1$, then
$\sum_{t=1}^{T}\Biggl(\sum_{j=0}^{t-1}q^{j}a_{t-j-1}\Biggr)^{2}\leq\frac{1}{(1-q)^{2}}\sum_{k=0}^{T-1}a_{k}^{2}.$
To show this, we expand the square and re-index with k=tj1k=t-j-1 and $m=t-\ell-1$, which yields
$\sum_{t=1}^{T}\Biggl(\sum_{j=0}^{t-1}q^{j}a_{t-j-1}\Biggr)^{2}=\sum_{k=0}^{T-1}\sum_{m=0}^{T-1}a_{k}a_{m}\sum_{t=\max\{k,m\}+1}^{T}q^{2t-k-m-2}.$
For $0<q<1$ we have
$\sum_{t=\max\{k,m\}+1}^{T}q^{2t-k-m-2}=q^{2\max\{k,m\}-k-m}\cdot\sum_{r=0}^{T-\max\{k,m\}-1}q^{2r}\leq\frac{q^{2\max\{k,m\}-k-m}}{1-q^{2}}.$
Hence
$\sum_{t=1}^{T}\Biggl(\sum_{j=0}^{t-1}q^{j}a_{t-j-1}\Biggr)^{2}\leq\frac{1}{1-q^{2}}\sum_{k=0}^{T-1}\sum_{m=0}^{T-1}q^{2\max\{k,m\}-k-m}a_{k}a_{m}.$
Observe that $q^{2\max\{k,m\}-k-m}=q^{|k-m|}$, so the right-hand side equals
$\frac{1}{1-q^{2}}\mathbf{a}^{\top}K\mathbf{a},\qquad K_{k,m}=q^{|k-m|},\qquad\mathbf{a}=(a_{0},\dots,a_{T-1})^{\top}.$
By the Schur test,
$\|K\|_{2}\leq\max_{k}\sum_{m=0}^{T-1}q^{|k-m|}\leq 1+2\sum_{r=1}^{\infty}q^{r}=\frac{1+q}{1-q}.$
Therefore,
$\sum_{t=1}^{T}\Biggl(\sum_{j=0}^{t-1}q^{j}a_{t-j-1}\Biggr)^{2}\leq\frac{1}{1-q^{2}}\|K\|_{2}\|\mathbf{a}\|_{2}^{2}\leq\frac{1}{1-q^{2}}\cdot\frac{1+q}{1-q}\sum_{k=0}^{T-1}a_{k}^{2}=\frac{1}{(1-q)^{2}}\sum_{k=0}^{T-1}a_{k}^{2},$
since $1-q^{2}=(1-q)(1+q)$. This proves the claim.
Appendix E Proof of Lemma C.2
Recall from (8) that
$V_{j}^{(t)}=\sum_{s=1}^{t}\mathbf{p}^{(s)}\mathbf{p}^{(s)\top}+\frac{\lambda}{c_{\mu_{j}}}I_{N},$
First, note:
$(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})=(p_{j}^{(t)},\mathbf{p}_{-j}^{(t)})+(\underbrace{p_{j}^{(t+1)}-p_{j}^{(t)}}_{\Delta p_{j}^{(t)}},\boldsymbol{0}_{-j})=\mathbf{p}^{(t)}+\Delta p_{j}^{(t)}\mathbf{e}_{j},$
where $\mathbf{e}_{j}$ is the null vector with only entry $1$ in the $j$-th coordinate. So:
$\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}\leq\|\mathbf{p}^{(t)}\|_{V_{j}^{(t),-1}}+\|\Delta p_{j}^{(t)}\mathbf{e}_{j}\|_{V_{j}^{(t),-1}}\leq\|\mathbf{p}^{(t)}\|_{V_{j}^{(t),-1}}+2\bar{p}_{j}\sqrt{[V_{j}^{(t),-1}]_{jj}}.$
Hence
$\sum_{t=0}^{T-1}\sum_{j\in[N]}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}\leq\underset{=A}{\underbrace{\sum_{j\in[N]}\sum_{t=0}^{T-1}\|\mathbf{p}^{(t)}\|_{V_{j}^{(t),-1}}}}+2\bar{p}\underset{=B}{\underbrace{\sum_{j\in[N]}\sum_{t=0}^{T-1}\sqrt{[V_{j}^{(t),-1}]_{jj}}}},$
where $\bar{p}=\max_{j\in[N]}\{\bar{p}_{j}\}$.
Bound B.
Fix $j\in[N]$. Let $\underline{p}=\min_{j\in[N]}\underline{p}_{j}>0$. Then for every $\mathbf{p}\in\mathcal{P}$ we have $\mathbf{p}\geq\underline{p}\mathbf{1}$ element-wise. For $t\geq 1$ and any $\mathbf{x}\in\mathbb{R}^{N}$,
$\mathbf{x}^{\top}\left(\sum_{s=1}^{t}\mathbf{p}^{(s)}\mathbf{p}^{(s)\top}\right)\mathbf{x}=\sum_{s=1}^{t}(\mathbf{x}^{\top}\mathbf{p}^{(s)})^{2}\geq\sum_{s=1}^{t}(\underline{p}x_{j})^{2}=t\underline{p}^{2}x_{j}^{2}=\mathbf{x}^{\top}(t\underline{p}^{2}\mathbf{e}_{j}\mathbf{e}_{j}^{\top})\mathbf{x},$
where $\mathbf{e}_{j}$ is the $j$-th vector of the canonical euclidean basis in $\mathbb{R}^{N}$. For simplicity of notation, let $\lambda_{j}=\frac{\lambda}{c_{\mu_{j}}}$. We have
$\sum_{s=1}^{t}\mathbf{p}^{(s)}\mathbf{p}^{(s)\top}\succeq t\underline{p}^{2}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\quad\Rightarrow\quad V_{j}^{(t)}=\lambda_{j}I_{N}+\sum_{s=1}^{t}\mathbf{p}^{(s)}\mathbf{p}^{(s)\top}\succeq\lambda_{j}I_{N}+t\underline{p}^{2}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}.$
Hence
$V_{j}^{(t),-1}\preceq\left(\lambda_{j}I_{N}+t\underline{p}^{2}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\right)^{-1}.$
And the $(j,j)$ element satisfies
$V_{j}^{(t),-1}]_{jj}=\mathbf{e}_{j}^{\top}V_{j}^{(t),-1}\mathbf{e}_{j}\leq\mathbf{e}_{j}^{\top}\left(\lambda_{j}I_{N}+t\underline{p}^{2}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\right)^{-1}\mathbf{e}_{j}=\frac{1}{\lambda_{j}+t\underline{p}^{2}},\implies\sqrt{[V_{j}^{(t),-1}]_{jj}}\leq\frac{1}{\sqrt{\lambda_{j}+t\underline{p}^{2}}}$
Because $x\mapsto(\lambda_{j}+\underline{p}^{2}x)^{-1/2}$ is decreasing on $0,\infty)$,
$\sum_{t=0}^{T-1}[V_{j}^{(t),-1}]_{jj}^{1/2}\leq\frac{1}{\sqrt{\lambda_{j}}}+\int_{0}^{T}\frac{dx}{\sqrt{\lambda_{j}+\underline{p}^{2}x}}=\frac{1}{\sqrt{\lambda_{j}}}+\frac{2}{\underline{p}^{2}}\left(\sqrt{\lambda_{j}+\underline{p}^{2}T}-\sqrt{\lambda_{j}}\right).$
Summing over $j\in[N]$ and retaining the dependence on $\lambda,N$ and $T$ we obtain $B=O\left(\frac{N}{\sqrt{\lambda}}+N\sqrt{\lambda+T}\right)$.
Bound A.
We use the Elliptical Potential Lemma in Carpentier et al. (2020) Proposition 1 (by replacing $p\leftarrow 1$ and $\lambda\leftarrow\frac{\lambda}{c_{\mu_{j}}}$):
$\sum_{j\in[N]}\sum_{t=0}^{T-1}\|\mathbf{p}^{(t)}\|_{V_{j}^{(t),-1}}$ $\leq\sum_{j\in[N]}\sqrt{TN\log\left(\frac{c_{\mu_{j}}T}{N\lambda}+1\right)}\leq N\sqrt{TN\log\left(T\frac{c_{\mu}}{N\lambda}+1\right)}.$
Merging the two bounds.
Retaining only the dependence on $\lambda,N$ and $T$, we recover
$\sum_{t=0}^{T-1}\sum_{j\in[N]}\|(p_{j}^{(t+1)},\mathbf{p}_{-j}^{(t)})\|_{V_{j}^{(t),-1}}=O\left(\frac{N}{\sqrt{\lambda}}+N\sqrt{\lambda+T}+N\sqrt{TN\log\left(\frac{T}{N\lambda}+1\right)}\right).$
This completes the proof.
Appendix F Proof of Lemma 2.6
We want to prove that, under 2.5, for every $i\in[N]$, $\eta_{i}^{(t)}$ is $L_{\mu_{i}}$-subgaussian conditionally on $\mathcal{H}_{i}^{(t-1)}$. Fix $i\in[N]$. We first prove that $\mathbb{E}[\eta_{i}^{(t)}|\mathcal{H}^{(t-1)}_{i}]=0$, where
$\eta_{i}^{(t)}=y_{i}^{(t)}-\mu_{i}(\langle\mathbf{p}^{(t)},\boldsymbol{\theta}_{i,0}\rangle).$
We recall that $p_{i}^{(t)}$ is computed using information $\mathcal{H}_{i}^{(t-1)}=\{(y_{i}^{(s)},\mathbf{p}^{(s)})\}_{s\leq t-1}$. Let
$\mathcal{F}_{i}^{(t-1)}=\sigma\left(\mathcal{H}^{(t-1)}_{i},\mathbf{p}^{(t)}\right).$
Since $\mathcal{H}_{i}^{(t-1)}\subseteq\mathcal{F}_{i}^{(t-1)}$, we have by the law of total expectation
$\mathbb{E}[\eta_{i}^{(t)}|\mathcal{H}^{(t-1)}_{i}$] $=\mathbb{E}[\mathbb{E}[\eta_{i}^{(t)}|\mathcal{F}^{(t-1)}_{i}]|\mathcal{H}^{(t-1)}_{i}$]
$=\mathbb{E}[\mathbb{E}[y_{i}^{(t)}-\mu_{i}(\langle\mathbf{p}^{(t)},\boldsymbol{\theta}_{i,0}\rangle)|\mathcal{F}^{(t-1)}_{i}]|\mathcal{H}^{(t-1)}_{i}$]
$=\mathbb{E}[\mu_{i}(\langle\mathbf{p}^{(t)},\boldsymbol{\theta}_{i,0}\rangle)-\mu_{i}(\langle\mathbf{p}^{(t)},\boldsymbol{\theta}_{i,0}\rangle)|\mathcal{H}^{(t-1)}_{i}$]
$=0.$
In the third equality we used that, by construction $y_{i}^{(t)}$ is independent of $\mathcal{H}_{i}^{(t-1)}$ given $\mathbf{p}^{(t)}$, i.e. $\mathbb{E}[y_{i}^{(t)}\mid\mathcal{H}_{i}^{(t-1)},\mathbf{p}^{(t)}]=\mathbb{E}[y_{i}^{(t)}\mid\mathbf{p}^{(t)}]=\mu_{i}(\langle\mathbf{p}^{(t)},\boldsymbol{\theta}_{i,0}\rangle)$. Now let $w_{i}^{(t)}\equiv\left\langle\boldsymbol{\theta}_{i,0},\mathbf{p}^{(t)}\right\rangle$ (this is $\mathcal{F}_{i}^{(t-1)}$-measurable), and recall from (2) that we model
$\frac{d\mathbb{P}_{\boldsymbol{\theta}_{i,0}}(y_{i}^{(t)}|\mathbf{p}^{(t)})}{d\nu_{i}(y_{i}^{(t)})}=\exp\left\{y_{i}^{(t)}w_{i}^{(t)}-b_{i}(w_{i}^{(t)})+c_{i}(y_{i}^{(t)})\right\},$
Then, with $\mu_{i}(w_{i}^{(t)})=b_{i}^{\prime}(w_{i}^{(t)})$ and $\eta_{i}^{(t)}=y_{i}^{(t)}-\mu_{i}(w_{i}^{(t)})$,
$\mathbb{E}\left[e^{\lambda\eta_{i}^{(t)}}|\mathcal{F}_{i}^{(t-1)}\right]$ $=e^{-\lambda\mu_{i}(w_{i}^{(t)})}\int e^{\lambda y_{i}^{(t)}}\exp\left\{y_{i}^{(t)}w_{i}^{(t)}-b_{i}(w_{i}^{(t)})+c_{i}(y_{i}^{(t)})\right\}d\nu_{i}(y_{i}^{(t)})$
$=e^{-\lambda\mu_{i}(w_{i}^{(t)})}e^{-b_{i}(w_{i}^{(t)})}\int\exp\left\{y_{i}^{(t)}(w_{i}^{(t)}+\lambda)+c_{i}(y_{i}^{(t)})\right\}d\nu_{i}(y_{i}^{(t)})$
$=e^{-\lambda\mu_{i}(w_{i}^{(t)})}e^{-b_{i}(w_{i}^{(t)})}e^{b_{i}(w_{i}^{(t)}+\lambda)}=\exp\left\{b_{i}(w_{i}^{(t)}+\lambda)-b_{i}(w_{i}^{(t)})-b_{i}^{\prime}(w_{i}^{(t)})\lambda\right\}.$
Taking logs gives the identity
$\log\mathbb{E}\left[\exp\big(\lambda\eta_{i}^{(t)}\big)\middle|\mathcal{F}_{i}^{(t-1)}\right]=b_{i}(w_{i}^{(t)}+\lambda)-b_{i}(w_{i}^{(t)})-b_{i}^{\prime}(w_{i}^{(t)})\lambda.$
By Taylor’s expansion, there exists a $\xi\in[0,1]$ such that
$b_{i}(w_{i}^{(t)}+\lambda)-b_{i}(w_{i}^{(t)})-b_{i}^{\prime}(w_{i}^{(t)})\lambda=\tfrac{1}{2}b_{i}^{\prime\prime}(w_{i}^{(t)}+\xi\lambda)\lambda^{2}=\tfrac{1}{2}\mu_{i}^{\prime}(w_{i}^{(t)}+\xi\lambda)\lambda^{2}.$
Assumption 2.5 yields $\mu_{i}^{\prime}(\cdot)\leq L_{\mu_{i}}$ on $\mathbb{R}$, hence
$\log\mathbb{E}\left[\exp\big(\lambda\eta_{i}^{(t)}\big)\middle|\mathcal{F}_{i}^{(t-1)}\right]\leq\frac{L_{\mu_{i}}\lambda^{2}}{2}\quad\implies\quad\mathbb{E}\left[\exp\big(\lambda\eta_{i}^{(t)}\big)\middle|\mathcal{F}_{i}^{(t-1)}\right]\leq\exp\left(\frac{L_{\mu_{i}}\lambda^{2}}{2}\right).$
And consequently
$\mathbb{E}\left[\exp\big(\lambda\eta_{i}^{(t)}\big)\middle|\mathcal{H}_{i}^{(t-1)}\right]=\mathbb{E}\left[\mathbb{E}\left[\exp\big(\lambda\eta_{i}^{(t)}\big)\middle|\mathcal{F}_{i}^{(t-1)}\right]\middle|\mathcal{H}_{i}^{(t-1)}\right]\leq\exp\left(\tfrac{L_{\mu_{i}}\lambda^{2}}{2}\right)\quad\text{a.s. for all $\lambda\in\mathbb{R}$}.$
This proves that $\eta_{i}^{(t)}$ is conditionally sub-gaussian with variance proxy $L_{\mu_{i}}$. This completes the proof.